OpenStudy (anonymous):

9.01], [9.02] Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: •Does the graph open up or down? How do you know? (2 points) •Explain whether the graph has a maximum or minimum point. (2 points) •Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (anonymous):

All you have to do is create an equation and finding the vertex and x intercept. :_

OpenStudy (anonymous):

Here's one for you: y = x^2 - 8x + 12 This factors as y = (x - 6)(x - 2) So, x-intercepts are at (6, 0), and (2, 0) This opens up because the implied "a" value is 1. It also has the vertex form of y = (x - 4)^2 - 4 so it has a vertex at (4, -4) and it is at this point where the graph and equation has its low-point.

OpenStudy (anonymous):

OMG! thanks so much! You r amazing!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

Now that's a nice smile!

OpenStudy (anonymous):

Thx!